the barber's profit "p" each week depends on his charge "c" per haircut. it is modeled by the equation p= -200c^2 + 2400c - 4700. what price should he charge for the largest profit? what would be his maximum profit?
Answer by nerdybill(7384)

the barber's profit "p" each week depends on his charge "c" per haircut. it is modeled by the equation p= -200c^2 + 2400c - 4700. what price should he charge for the largest profit? what would be his maximum profit?
.
By examining the given equation (coefficient associated with the c^2 term), we know that it is a parabola that opens downwards. Therefore, the vertex is the max.
.
what price should he charge for the largest profit?
c = -b/(2a)
c = -2400/(2(-200))
c = -2400/(-400)
c = $6
.
what would be his maximum profit?
plug the value above back into:
p= -200c^2 + 2400c - 4700
p= -200(6)^2 + 2400(6) - 4700
p= -200c^2 + 2400c - 4700
p = $2500